# Optimal e-value testing for properly constrained hypotheses

### v1

@@ Abstract
Hypothesis testing via e-variables can be framed as a sequential betting game, where a player each round picks an e-variable. A good player’s strategy results in an effective statistical test that rejects the null hypothesis as soon as sufficient evidence arises. Building on recent advances, we address the question of restricting the pool of e-variables to simplify strategy design without compromising effectiveness. We extend the results of Clerico (2024), by characterising optimal sets of e-variables for a broad class of non-parametric hypothesis tests, defined by finitely many regular constraints. As an application, we discuss optimality in algorithmic mean estimation, including the case of heavy-tailed random variables.
@@ 1 Introduction
Hypothesis testing is the branch of statistics concerned with verifying if an observed data set is consistent with some given theory, typically referred to as the null hypothesis. Traditionally, this is formulated in terms of p-values, which represent the probability of obtaining an empirical test statistic at least as “extreme” as the one observed, under the assumption that the null hypothesis holds. However, most classical p-value methods are designed for single, fixed-sample experiments. In modern research, where data are often collected iteratively or experiments are repeated, recalculating p-values with additional data can lead to false positives and misleading conclusions. To overcome these limitations, recent developments propose using e-values as a more reliable alternative for sequential and adaptive hypothesis testing.
@@
E-values quantify evidence against a hypothesis and allow flexible data collection strategies, such as optional stopping or continuation, and post hoc significance level tuning. While related ideas were developed decades ago by early works (Wald, 1945; Darling and Robbins, 1967; Lai, 1976; Siegmund, 1978), interest in e-values surged with a series of papers whose first preprints appeared over the space of few months in 2019 (Wasserman et al., 2020; Shafer, 2021; Vovk and Wang, 2021; Grünwald et al., 2024). Hypothesis testing with e-values can be phrased in terms of a sequential betting game (Ramdas et al., 2023), adopting a game-theoretic perspective on probability recently developed by Shafer and Vovk (2001, 2019). In this game, at each round the player selects an e-variable, namely a non-negative function with an expected value of at most one under the null hypothesis. The player then receives a reward, determined by the value of the picked e-variable at the next observed data point. If a significant grow of the cumulative reward is observed, the null hypothesis can be confidently rejected, as standard martingale analysis shows that this is unlikely to happen under the null hypothesis. We refer to Ramdas and Wang (2024) for an overview on e-value testing and game-theoretic interpretations.
@@
A testing approach is effective if it leads to the rejection of a hypothesis as soon as there is enough evidence against it in the data. In our framework, this can be achieved if the player adopts a sound strategy, which yields high rewards whenever possible. The theoretical foundation for selecting such strategies in sequential games lies in online learning theory (Cesa-Bianchi and Lugosi, 2006; Orabona, 2023). To simplify strategy design and leverage online learning theoretical tools, it is often convenient to restrict the pool of e-variables among which the player can pick each round. In order to make this point more explicit, Clerico (2024) introduced the concept of restricted testing game. A key question is whether such restriction is detrimental for the test, by preventing the selection of strategies that could yield higher rewards. The answer is no, provided the restriction is well chosen. Moreover, in some cases, there exists an optimal class of e-variables to which one should restrict the game, informally the smallest set within which any good strategy must choose the e-variables. These ideas were developed in Clerico (2024), in the context of a simple e-variable testing problem that was motivated by mean estimation for a random variable bounded in $[0,1]$. This work extends these results to a wider class of hypothesis testing problems, by characterising the optimal set of e-variables for non-parametric hypotheses defined by a finite number of suitably regular constraints. This class is broad enough to capture meaningful problems, while allowing us to use standard analysis techniques, without relying on advanced tools from functional analysis and measure theory.
Notions of optimality for single e-variables, such as log-optimality (Koolen and Daniel, 2022; Larsson et al., 2024; Grünwald et al., 2024) and admissibility (Ramdas et al., 2022), have been previously extensively explored in the sequential hypothesis testing literature. However, as highlighted by Clerico (2024), the key distinction in our approach is the emphasis on optimality at the level of sets of e-variables, rather than focusing on individual e-variables.
@@ 1.1 Structure
This paper is organised as follows. In Section 2, we briefly review the fundamental standard concepts related to e-value hypothesis testing. Section 3 introduces the key definitions and basic properties of majorising and optimal classes of e-variables, while Section 4 defines the concept of properly constrained hypothesis, which serves as the central framework for this work. Section 5 discusses the dual class of e-variables for properly constrained hypotheses and states our main result: the optimality of the dual class (Theorem 1), which will be proved throughout the subsequent sections. Specifically, we begin by addressing the case of a finite sample space in Section 6, and then introduce the notion of matching sets in Section 7, which will be an essential tool to extend the optimality results to general closed sample spaces in Section 8. In Section 9, we broaden our results, by addressing tests of finitely (but not properly) constrained hypotheses, and by considering hypotheses defined via inequality constraints. In Section 10, our results are applied to the problem of deriving confidence sequences for the mean of (bounded and heavy-tailed) random variables. Lastly, Section 11 provides concluding remarks and outlines potential directions for future research.
@@ 1.2 Notation
Sequences are denoted as $(x_{t})_{t\geq T_{0}}$, where $t$ is an integer-valued index and $T_{0}$ its smallest value. For high-probability statements, $\mathbb{P}$ expresses probability with respect to all the involved randomness. For instance, if $(X_{t})_{t\geq 1}$ is a sequence of independent draws from $P\in\mathcal{P}_{\mathcal{X}}$, we can write $\mathbb{P}\big{(}X_{t}\geq 0\,,\,\forall t\geq 1\big{)}$, with obvious meaning.
@@ 2 Algorithmic hypothesis testing
  • measurably [^2] We assume that the choice of $E_{t}$ is measurable, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable. picks an e-variable $E_{t}\in\mathcal{E}$, based solely on the past observations $x_{1},\dots,x_{t-1}$;
@@
The cumulative reward of the player can be interpreted as evidence accumulated against $\mathcal{H}$. Specifically, Lemma 1 justifies a sequential testing procedure, where the null hypothesis (asserting that the data are generated from a $P\in\mathcal{H}$) is rejected whenever the player’s cumulative reward exceeds the threshold value of $\log(1/\delta)$. Here, $\delta\in(0,1)$ denotes the type II confidence level, meaning that if the null hypothesis is true, it will be rejected with probability at most $\delta$. It is worth noticing that the high-probability inequality in Lemma 1 holds uniformly for any data set size $T$. This ensures that when data are collected sequentially, optional stopping or continuation are allowed, an advantage not afforded by standard p-value methods. For a comprehensive overview of sequential testing by betting, we refer to Chapter 6 of Ramdas and Wang (2024).
@@ 3 Majorising e-classes and optimal e-class
We now present the definitions and basic properties of majorising and optimal e-classes, as introduced in Clerico (2024). In the following, fix a closed $\mathcal{X}\subseteq\mathbb{R}^{n}$ and a hypothesis $\mathcal{H}$ on $\mathcal{X}$. We consider the standard partial ordering on the set of real functions on $\mathcal{X}$. Specifically, given two functions $f$ and $f^{\prime}$ from $\mathcal{X}$ to $\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in\mathcal{X}$. If $f\succeq f^{\prime}$ and $f\neq f^{\prime}$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$. We will also use the symbols $\preceq$ and $\prec$, with obvious meaning.
@@
By Lemma 2, it is enough to show that the set of all maximal e-variables is a majorising e-class, or equivalently that every e-variable is majorised by a maximal e-variable. So, fix $E_{0}\in\mathcal{E}_{\mathcal{H}}$. The set $\mathcal{E}_{0}=\{E\in\mathcal{E}_{\mathcal{H}}\,:\,E\succeq E_{0}\}\subseteq\mathcal{E}_{\mathcal{H}}$ is closed, hence compact. Moreover, the mapping $E\mapsto E(x_{1})$ is continuous on $\mathcal{E}_{\mathcal{H}}$. We can thus find $E_{1}\in\operatorname*{arg\,max}_{E\in\mathcal{E}_{0}}E(x_{1})$. We iterate this procedure up to $i=d$, with $\mathcal{E}_{i}=\{E\in\mathcal{E}_{\mathcal{H}}\,:\,E\succeq E_{i}\}$ and $E_{i}\in\operatorname*{arg\,max}_{E\in\mathcal{E}_{i-1}}E(x_{i})$. Now, by construction $E_{d}\succeq E_{0}$. Consider any e-variable $E^{\star}\succeq E_{d}$ and fix any integer index $i\in[1,d]$. Then, $E^{\star}\succeq E_{d}\succeq E_{i-1}$, which by the definition of $E_{i}$ implies that $E^{\star}(x_{i})\leq E_{i}(x_{i})\leq E_{d}(x_{i})$, and so $E^{\star}(x_{i})=E_{d}(x_{i})$. Therefore, $E^{\star}=E_{d}$ and $E_{d}$ is maximal. ∎
@@
One might think that if $\sup_{P\in\mathcal{H}}P(\{x\})>0$ for every $x\in\mathcal{X}$ the optimal e-class always exists. However, this is not always the case and a counter example is provided in Appendix C.
@@ 4 Properly constrained hypotheses
In the remainder of this paper, we will mainly focus on a specific kind of hypotheses, for which we aim to fully characterise the optimal e-class. Notably, Clerico (2024) examined the simple case where $\mathcal{X}=[0,1]$ and $\mathcal{H}$ is the set of probability measures on $\mathcal{X}$ with a fixed mean. Here, we discuss the more general setting where $\mathcal{X}$ is a generic closed set in $\mathbb{R}^{n}$ and the hypothesis $\mathcal{H}$ is defined in terms of the expectation of a continuous function $\Phi:\mathcal{X}\to\mathbb{R}^{m}$. Since the approach in Clerico (2024) is specifically designed for their simpler setting, we will have to develop more sophisticated proof techniques to handle the broader framework that we consider in this work.
@@
We remark that any finitely constrained $\mathcal{H}$ is convex. Indeed, $\mathcal{P}_{\Phi}$ is convex, and $\mathcal{H}$ is a level set of the convex functional $P\mapsto\langle P,\Phi\rangle$ on $\mathcal{P}_{\Phi}$. Next, we list a few basic properties of properly constrained hypotheses. The proofs, detailed in Appendix A, are generally rather short and rely on elementary tools from linear algebra and convex analysis.
@@
Lemma 6.
Lemma 7.
Lemma 8.
@@
As a side note, we state a stronger version of the last claim in Lemma 8, which may be of future interest, while not directly used in this work. The proof is deferred to Appendix D.
Proposition 1.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. There is a continuous function $A:\mathcal{X}\to[1,+\infty)$ such that $\langle P,A\rangle<+\infty$ for every $P\in\mathcal{H}$, $\sup_{P\in\mathcal{H}}P(\{x\})\geq 1/A(x)$ for all $x\in\mathcal{X}$, and $E\preceq A$ for any $E\in\mathcal{E}_{\mathcal{H}}$.

### v3

@@ Abstract
Hypothesis testing via e-variables can be framed as a sequential betting game, where a player each round picks an e-variable. A good player’s strategy results in an effective statistical test that rejects the null hypothesis as soon as sufficient evidence arises. Building on recent advances, we address the question of restricting the pool of e-variables to simplify strategy design without compromising effectiveness. We extend the results of Clerico (2024), by characterising optimal sets of e-variables for a broad class of non-parametric hypothesis tests, defined by finitely many regular constraints. As an application, we discuss this notion of optimality in algorithmic mean estimation, including for heavy-tailed random variables.
@@ 1 Introduction
Hypothesis testing is the branch of statistics concerned with verifying if an observed data set is consistent with a given theory, often referred to as the null hypothesis. Traditionally, this is formulated in terms of p-values, which represent the probability of obtaining an empirical test statistic at least as “extreme” as the one observed, assuming that the null hypothesis holds. However, most classical p-value methods are designed for single, fixed-sample experiments. In modern research, where data are often collected iteratively or experiments are repeated, recalculating p-values with additional data can lead to false positives and misleading conclusions. To overcome these limitations, recent developments propose using e-values as a more reliable alternative for sequential and adaptive hypothesis testing.
@@
E-values quantify evidence against a hypothesis and allow flexible data collection strategies, such as optional stopping or continuation, and post hoc significance level tuning. While related ideas were developed decades ago by early works (Wald, 1945; Darling and Robbins, 1967; Lai, 1976; Siegmund, 1978), interest in e-values surged with a series of papers whose first preprints appeared over the space of few months in 2019 (Wasserman et al., 2020; Shafer, 2021; Vovk and Wang, 2021; Grünwald et al., 2024). In the present work, we focus on a particular case of testing with e-values, which can be phrased in terms of a sequential betting game (Ramdas et al., 2023), adopting a game-theoretic perspective on probability recently developed by Shafer and Vovk (2001, 2019). In this game, at each round the player selects a single round e-variable, namely a non-negative function with an expected value of at most one under the null hypothesis. The player then receives a reward, determined by the value of the picked e-variable at the next observed data point. If a significant grow of the cumulative reward is observed, the null hypothesis can be confidently rejected, as standard martingale analysis shows that this is unlikely to happen under the null hypothesis. We refer to Ramdas and Wang (2024) for an overview on e-value testing and game-theoretic interpretations.
@@
A testing approach is effective if it leads to the rejection of a hypothesis as soon as there is enough evidence against it in the data. In our framework, this can be achieved if the player adopts a sound strategy, which yields high rewards whenever possible. The theoretical foundation for selecting such strategies in sequential games lies in online learning theory (Cesa-Bianchi and Lugosi, 2006; Orabona, 2023). To simplify strategy design and leverage online learning theoretical tools, it is often convenient to restrict the pool of e-variables among which the player can pick each round. In order to make this point more explicit, Clerico (2024) introduced the concept of restricted testing game. A key question is whether such restriction is detrimental for the test, by preventing the selection of strategies that could yield higher rewards. Tor many instantiations of ths testing by betting framework the answer is no, provided the restriction is well chosen. Moreover, in some cases, there exists an optimal class of e-variables to which one should restrict the game, informally the smallest set within which any good strategy must choose the e-variables. These ideas were developed in Clerico (2024), in the context of a simple e-variable testing problem that was motivated by mean estimation for a random variable bounded in $[0,1]$. This work extends these results to a wider class of hypothesis testing problems, by characterising the optimal set of single round e-variables for non-parametric hypotheses defined by a finite number of suitably regular constraints. This class is broad enough to capture meaningful problems, while allowing us to use standard analysis techniques, without relying on advanced tools from functional analysis and measure theory.
Notions of optimality for single e-variables, such as log-optimality (Koolen and Daniel, 2022; Larsson et al., 2024; Grünwald et al., 2024) and admissibility (Ramdas et al., 2022), have been previously extensively explored in the sequential hypothesis testing literature. However, as highlighted by Clerico (2024), the key distinction in our approach is the emphasis on optimality at the level of sets of e-variables, rather than focusing on individual e-variables. More explicitly, we show that for a class of constrained hypotheses, the whole set of single-round e-variables is dominated by the set of all maximal e-variables, namely e-variables that cannot be dominated by any other e-variables. We remark here that the notion of maximality of an e-variable is essentially equivalent to that of admissibility in classical statistics. Somehow our result can be seen as an existence and characterisation theorem of a minimal complete class of admissible e-variables. However, in the e-variable literature, it is common to define admissibility with respect to an alternative hypothesis (see Ramdas et al., 2022), while the current paper does not rely on any given alternative. For this reason we prefer to keep the term maximal (typical in the literature on posets), as done in Clerico (2024).
As a final remark, we stress once more that the results of this paper only focus on optimality of testing by betting via e-variables. This is by far not the most general type of sequential testing. An important open question is to characterise optimality for more general forms of sequential testing. More concretely, this paper is only one first step towards characterising all the admissible sequential tests, as it characterises the form of maximal single-round e-variables.
@@ 1.1 Structure
This paper is organised as follows. In Section 2, we briefly review e-value hypothesis testing by betting. Section 3 introduces the key definitions and basic properties of majorising and optimal classes of e-variables, while Section 4 defines the concept of properly constrained hypothesis, which serves as the central framework for this work. Section 5 discusses the dual class of e-variables for properly constrained hypotheses and states our main result: the optimality of the dual class (Theorem 1), which will be proved throughout the subsequent sections. Specifically, we begin by addressing the case of a finite sample space in Section 6, and then introduce the notion of matching sets in Section 7, which will be an essential tool to extend the optimality results to general closed sample spaces in Section 8. In Section 9, we broaden our results, by addressing tests of finitely (but not properly) constrained hypotheses, and by considering hypotheses defined via inequality constraints. In Section 10, our results are applied to the problem of deriving confidence sequences for the mean of (bounded and heavy-tailed) random variables. Lastly, Section 11 provides concluding remarks and outlines potential directions for future research.
@@ 1.2 Notation
Sequences are denoted as $(x_{t})_{t\geq T_{0}}$, where $t$ is an integer-valued index and $T_{0}$ its smallest value. For high-probability statements, $\mathbb{P}$ expresses probability with respect to all the involved randomness. For instance, if $(X_{t})_{t\geq 1}$ is a sequence of independent draws from $P\in\mathcal{P}_{\mathcal{X}}$, we write $\mathbb{P}\big{(}X_{t}\geq 0\,,\,\forall t\geq 1\big{)}$.
@@ 2 Algorithmic hypothesis testing
We remark that we are here focused only on single-round hypotheses and e-variables. In particular, we are not considering hypotheses on a whole sequence of observations (in $\mathcal{X}^{\infty}$, as done for instance in Ramdas et al., 2022), but only on $\mathcal{X}$. In a way, we will be given for granted that the data come from as independent draws from some distribution on $\mathcal{X}$.
@@
  • measurably [^2] We assume that the choice of $E_{t}$ is measurable, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable. selects $E_{t}\in\mathcal{E}$, based solely on the previous observations $x_{1},\dots,x_{t-1}$;
@@
The cumulative reward of the player can be interpreted as evidence accumulated against $\mathcal{H}$. Specifically, Proposition 1 justifies a sequential testing procedure, where the null hypothesis (asserting that the data are generated from a $P\in\mathcal{H}$) is rejected whenever the player’s cumulative reward exceeds the threshold value of $\log(1/\delta)$. Here, $\delta\in(0,1)$ denotes the type I confidence level, meaning that if the null hypothesis is true, it will be rejected with probability at most $\delta$. It is worth noticing that the high-probability inequality in Proposition 1 holds uniformly for any data set size $T$. This ensures that when data are collected sequentially, optional stopping or continuation are allowed, an advantage not afforded by standard p-value methods. For a comprehensive overview of sequential testing by betting, we refer to Chapter 6 of Ramdas and Wang (2024).
@@ 3 Majorising e-classes and optimal e-class
We now present the definitions and basic properties of majorising and optimal e-classes, as introduced in Clerico (2024). In the following, fix a closed $\mathcal{X}\subseteq\mathbb{R}^{n}$ and a hypothesis $\mathcal{H}$ on $\mathcal{X}$. We consider the standard partial ordering for the real functions on $\mathcal{X}$: given two functions $f$ and $f^{\prime}$ from $\mathcal{X}$ to $\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in\mathcal{X}$. If $f\succeq f^{\prime}$ and $f\neq f^{\prime}$, $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$. We will also use the symbols $\preceq$ and $\prec$, with obvious meaning.
@@
By Lemma 1, it is enough to show that the set of all maximal e-variables is a majorising e-class, or equivalently that every e-variable is majorised by a maximal e-variable. So, fix $E_{0}\in\mathcal{E}_{\mathcal{H}}$. The set $\mathcal{E}_{0}=\{E\in\mathcal{E}_{\mathcal{H}}\,:\,E\succeq E_{0}\}\subseteq\mathcal{E}_{\mathcal{H}}$ is closed, hence compact. Moreover, the mapping $E\mapsto E(x_{1})$ is continuous on $\mathcal{E}_{\mathcal{H}}$. We can thus find $E_{1}\in\operatorname*{arg\,max}_{E\in\mathcal{E}_{0}}E(x_{1})$. We iterate this procedure up to $i=d$, with $\mathcal{E}_{i}=\{E\in\mathcal{E}_{\mathcal{H}}\,:\,E\succeq E_{i}\}$ and $E_{i}\in\operatorname*{arg\,max}_{E\in\mathcal{E}_{i-1}}E(x_{i})$. By construction, $E_{d}\succeq E_{0}$. Consider any e-variable $E^{\star}\succeq E_{d}$ and fix any integer index $i\in[1,d]$. Then, $E^{\star}\succeq E_{d}\succeq E_{i-1}$, which by the definition of $E_{i}$ implies that $E^{\star}(x_{i})\leq E_{i}(x_{i})\leq E_{d}(x_{i})$, and so $E^{\star}(x_{i})=E_{d}(x_{i})$. Thus, $E^{\star}=E_{d}$ and $E_{d}$ is maximal. ∎
@@
One might think that if $\sup_{P\in\mathcal{H}}P(\{x\})>0$ for every $x\in\mathcal{X}$ the optimal e-class always exists. However, this is not always the case, as shown by the next result (whose proof is deferred to Appendix A). Remarkably, this shows that allowing the e-variables to take the value +$\infty$+ ∞ would not be enough to guarantee the existence of the optimal e-class.
Proposition 2.
Let $\mathcal{X}=[0,1]$. Consider the hypothesis
=$\mathcal{H}=\big{\{}P\in\mathcal{P}_{[0,1]}\,:\,P(\{0\})\geq 1/2\big{\}}\cup\{U_{[0,1]}\}\,,$
with $U_{[0,1]}$ the uniform distribution on $[0,1]$. $\mathcal{H}$ does not admit an optimal e-class.
@@ 4 Properly constrained hypotheses
In the remainder of this paper, we will mainly focus on a specific kind of hypotheses, for which we aim to fully characterise the optimal e-class. Notably, Clerico (2024) examined the simple case where $\mathcal{X}=[0,1]$ and $\mathcal{H}$ is the set of probability measures on $\mathcal{X}$ with a fixed mean. More explicitly, for $\mu\in(0,1)$, define
$\mathcal{H}_{\mu}=\{P\in\mathcal{P}_{[0,1]}\,:\,\langle P,X\rangle=\mu\}$
. Testing for $\mathcal{H}_{\mu}$ means that we are testing if the mean of a bounded random variable is $\mu$. Clerico (2024) showed that the optimal e-class exists and is presisely
$\mathcal{E}_{\mu}=\big{\{}x\mapsto 1+\lambda(x-\mu)\,:\,\lambda\in[1/(\mu-1),1/\mu]\big{\}}$
(see Section 10.1 for more details).
We remark that the hypothesis $\mathcal{H}_{\mu}$ is an example of a constrained hypothesis, namely a hypothesis that can be defined via the linear constraint $\langle P,X-\mu\rangle=0$. Here, we discuss the existence and characterisation of the optimal e-class for hypothesis of a similar form. We consider the general setting where $\mathcal{X}$ is a generic closed set in $\mathbb{R}^{n}$ and the hypothesis $\mathcal{H}$ is defined in terms of the expectation of a continuous function $\Phi:\mathcal{X}\to\mathbb{R}^{m}$. As an application (see Section 10.2) we will for instance be able to characterise the optimal e-class for testing the mean of heavy tail random variables. Since the approach in Clerico (2024) is specifically designed for their simpler setting, we will have to develop more sophisticated proof techniques to handle the broader framework that we consider in this work.
@@
Next, we comment on a few properties of finitely and properly constrained hypotheses. All proofs (often short and building on elementary results from real algebra and convex analysis) are omitted here, and deferred to Appendix A.
Any finitely constrained $\mathcal{H}$ is convex, as $\mathcal{P}_{\Phi}$ is convex and $\mathcal{H}$ is a level set of the convex functional $P\mapsto\langle P,\Phi\rangle$ on $\mathcal{P}_{\Phi}$.
The reason for focusing on properly constrained hypotheses is that if 00 is on the (relative) boundary of $\mathcal{C}$, then one can show that all $P\in\mathcal{H}$ need to be supported on the relative boundary (see Lemma 23). In particular, Corollary 1 states that in such case the optimal e-class does not exist. However, in such cases one can restrict $\mathcal{X}$ to a smaller set, and obtain a proper constraint from a non-proper one (this will be later formalised in Lemma 18). On the other hand, if a hypothesis is properly constrained, it is reach enough to put mass on every $x$. This is ensured by the next lemma (proved in Appendix A).
Lemma 4.
Not only we can find a measure in $\mathcal{H}$ that put mass on every $x$ if $\mathcal{H}$ is properly constrained, we can also ask for such measure to have finite support, a property that will turn out to be useful for our analysis.
Lemma 5.
In general, proper constraints might be redundant, in the sense that their components might be linear dependent. When this is not the case, they are minimal constraints. This linear independence of the components of minimal constraints will be a key tool for our results, as it will often allow to reduce the problem to analysis of finite-dimensional spaces.
Lemma 6.
The above results suggest that, if we have a proper constraint, it can be transformed in a minimal constraint by eliminating some of its components. In particular, we expect that we can construct a minimal constraint for every properly constrained hypothesis. This is indeed the case.
@@
Last, we conclude with a characterisation of properly constrained hypotheses.